In triangle ABC, AC = 6cm and BC = 6 cm.
It is correct to affirm that either:
a) 4*Sin(A) = 3*Sin(C)
b) Sin(B) = Sin(C)
c) 3*Sin(B) = 4*Sin(C)
d) 6*Sen(A) = Sin(C)

Answer by ikleyn(52799)

.
In triangle ABC, AC = 6cm and BC = 6 cm.
It is correct to affirm that either:
a) 4*Sin(A) = 3*Sin(C)
b) Sin(B) = Sin(C)
c) 3*Sin(B) = 4*Sin(C)
d) 6*Sen(A) = Sin(C)
~~~~~~~~~~~~~~~~~~~~

ABC is isosceles triangle with the base side AB and lateral sides AC and BC of 6 cm each.
The angles A and B are the angles at the base.

a) 4*Sin(A) = 3*Sin(C)    - not necessary.
b) Sin(B) = Sin(C)        - True; Correct.
c) 3*Sin(B) = 4*Sin(C)    - not necessary.
d) 6*Sen(A) = Sin(C)      - not necessary.